OpenStudy (anonymous):

MEDAL AND FAN OF ANYONE WHO HELPS Choose the correct simplification of the expression (8x − 5)(2x2 − 5x − 6).

OpenStudy (anonymous):

medal and fan if you help!

OpenStudy (anonymous):

16x3 − 50x2 − 23x − 30 16x3 − 50x2 − 23x + 30 16x3 − 30x2 − 23x + 30 16x3 + 50x2 − 23x + 30

OpenStudy (johnweldon1993):

you multiply the first term in the first parenthesis by every term in the second parenthesis...then you multiply the second term in the first parenthesis by every term in the second parenthesis...and then simplify

OpenStudy (anonymous):

16x^3 - 40x^2 -48x - 10x^2 + 25x + 30 = 16x^3 - 50x^2 - 23x + 30

OpenStudy (anonymous):

thank you all so much!

OpenStudy (zpupster):

just the first term in the binomial by each term in the trinomial 8x x 2xe^2 = 16 x^3 8x x -5 x = -40x^2 8x x - 6 = -48 x now multiplying by the 2nd term in the binomail -5 x 2x^2 = -10x^2 -5 x - 5x =25x -5 x - 6 = 30 16^3-40x^2-10x^2 - 48 x+25 x +30 combine our like terms 16x^3 -50x^2-23x+30

OpenStudy (johnweldon1993):

no problem :)

OpenStudy (zpupster):

practice these a lot so they become 2nd nature . it will help you later